# Cherry picking in forests: A new characterization for the unrooted hybrid number of two phylogenetic trees

### v3

@@ 6. Bounding the weight of a cherry picking sequence from above by the hybrid number
Now, in case $r({\mathcal{N}})>0$, we essentially repeatedly look for non cut-edges $e$ in ${\mathcal{N}}$ that are contained in the image of one of the forests displayed by ${\mathcal{N}}$ but not in the other. If we can find such an edge $e$, by carefully considering how the images of the forests are embedded in ${\mathcal{N}}$, we can then complete the proof by removing $e$ from ${\mathcal{N}}$, applying Lemmas 5.4 and 5.5 and some (C2)-reduction, and using the fact that $r({\mathcal{N}})>r({\mathcal{N}}-e)$. We use this tool within a series of claims to show that either we can prove the theorem by induction, or that we can further restrict the way that the forests are embedded in ${\mathcal{N}}$. This eventually allows us to finally apply our induction hypothesis one final time and thus complete the proof.
More specifically, after first showing that we can assume that ${\mathcal{N}}$ must be pendantless (Claim 1), we then show that we can also assume that ${\mathcal{N}}$ must have a pendant blob ${\mathcal{B}}$ with at least two leaves such that no leaf in ${\mathcal{B}}$ is a component of ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$ and every edge in ${\mathcal{B}}$ is in ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ (Claims 2, 3 and 4). In Claims 5 and 6 we then show that either ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$ must contain a cherry $(x,y)$ that only satisfies Observation 3.1(ii). Then in Claims 7–11 we show that we can assume that there is a component in ${\mathcal{F}}$ that contains a cherry $(x,y)$ and a component in ${\mathcal{F}}^{\prime}$ that contains a cherry $(x,z)$ with $x,y,z\in L({\mathcal{B}})$. This allows us to assume that $((x,y),p)$ or $((x,y),(p,q))$ is a pendant subtree of one of the forests for some $x,y,p,q\in L({\mathcal{B}})$ distinct. This permits us to carry out a case analysis in Claims 12-14 in order to complete the proof.
@@
Base Case: Suppose $r({\mathcal{N}})=0$. Then ${\mathcal{N}}$ is a phylogenetic tree on $X$. If $|X|=1$, then $X=\{x\}$, some $x$, and $(x)$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with $w((x))=0$. Now, suppose $m=|X|>1$. Order the elements in $X$ as $x_{1},x_{2},\dots,x_{m}$ so that, for all $11\leq i\leq m-$ 1, $x_{i}$ is contained in a cherry of ${\mathcal{N}}|{\{x_{i},\dots,x_{m}\}}$ which is possible since any phylogenetic tree with two or more leaves always contains a cherry. Then, by applying Lemma 6.2 and using either a (C1) or (C3) reduction for each $x_{i}$, $11\leq i\leq m-$ 1, depending on whether (a) or (b) holds for $x_{i}$, respectively, it follows that $(x_{1},\dots,x_{m})$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with $r({\mathcal{N}})=w((x_{1},\dots,x_{m}))=0$. Hence the base case holds.
@@
Note first that we can and will assume from now on that $|X|\geq 3$ since for $|X|=1,2$, we can clearly find a cherry picking sequence $\sigma$ such that $w(\sigma)=0$ and so $r({\mathcal{N}})>0=w(\sigma)$ and hence (CP) holds.
Claim 0: We can and will assume that there is no $x\in X$ such that ${\mathcal{N}}-x$ is a phylogenetic network on $X\setminus\{x\}$ and either (a) there is a cherry $(x,y)$ contained in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, $y\in X\setminus\{x\}$, or (b) $x$ is a component in either ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$.
@@
Proof of Claim 0: Suppose that ${\mathcal{N}}$, ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ are such that Claim 0 does not hold, that is, there is some $x\in X$ such that ${\mathcal{N}}-x$ is a phylogenetic network on $X\setminus\{x\}$, and (a) or (b) holds. We first show that by applying Lemma 6.2 and (C1) and (C3)-reductions, we can transform ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ into forests ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ on a subset $X^{\prime}$ of $X$ and ${\mathcal{N}}$ into a phylogenetic network ${\mathcal{N}}^{\prime}$ on $X^{\prime}$ that displays ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ so that there is no $x\in X^{\prime}$ such that ${\mathcal{N}}^{\prime}-x$ is a phylogenetic network on $X^{\prime}\setminus\{x\}$, and (a) or (b) holds. Under the assumption that there exists a cherry picking sequence $\sigma^{\prime}$ for ${\mathcal{G}}^{\prime}$ and ${\mathcal{G}}$ with $r({\mathcal{N}}^{\prime})\geq w(\sigma^{\prime}),$ we then show that (CP) holds for ${\mathcal{N}}$. To see the first part, note that $|X\setminus\{x\}|\geq 2$ must hold as otherwise ${\mathcal{N}}-x$ would not be a phylogenetic network on $X\setminus\{x\}$. Now, starting with ${\mathcal{N}}$, repeatedly look for elements $y\in X$ that satisfy (a) or (b), and such that removing $y$ and $e_{y}$ and suppressing the resulting degree-2 vertex results in a phylogenetic network until this is no longer possible. Then we obtain a sequence $(x_{1},\dots,x_{m})$, $x_{i}\in X$, $m\geq 1$, where, for ${\mathcal{N}}_{1}={\mathcal{N}}$, ${\mathcal{N}}_{i+1}={\mathcal{N}}_{i}-x_{i}$, ${\mathcal{F}}[1]={\mathcal{F}}$, ${\mathcal{F}}[i+1]={\mathcal{F}}[i]-x_{i}$, ${\mathcal{F}}^{\prime}[1]={\mathcal{F}}^{\prime}$ and ${\mathcal{F}}^{\prime}[i+1]={\mathcal{F}}^{\prime}[i]-x_{i}$, $1\leq i\leq m$, we have for all $1\leq i\leq m$ that $x_{i}$ satisfies (a) or (b), and ${\mathcal{N}}_{i+1}={\mathcal{N}}_{i}-x_{i}$ is a phylogenetic network on $X\setminus\{x_{1},\dots,x_{i}\}$, ${\mathcal{N}}_{i}$ displays the forests ${\mathcal{F}}[i]$, ${\mathcal{F}}^{\prime}[i]$ and, by Lemma 6.2, $r({\mathcal{N}}_{i})=r({\mathcal{N}})$. In particular, ${\mathcal{N}}_{m+1}$ is a phylogenetic network on $X\setminus\{x_{1},\dots,x_{m}\}$ with $r({\mathcal{N}}_{{\color[rgb]{0,0,0}m+1}})=q$ that displays ${\mathcal{F}}[m+1]$ and ${\mathcal{F}}^{\prime}[m+1]$, and there is no $x\in X\setminus\{x_{1},\dots,x_{m}\}$ such that ${\mathcal{N}}_{m+1}-x$ is a phylogenetic network and either (a) or (b) holds.
@@
To complete the proof of the claim assume that we can show that (CP) holds for ${\mathcal{N}}_{m+1}$ along with the forests ${\mathcal{F}}[m+1]$ and ${\mathcal{F}}^{\prime}[m+1]$. Then there must exist a cherry picking sequence $\sigma^{\prime}$ for ${\mathcal{F}}[m+1]$ and ${\mathcal{F}}^{\prime}[m+1]$ with $r({\mathcal{N}}_{m+1})\geq w(\sigma^{\prime})$. Hence, $\sigma=(x_{1},\dots,x_{m})\circ\sigma^{\prime}$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. Since we have applied a (C1) or (C3) reduction to each $x_{i}$ depending on whether (a) or (b) holds for $x_{i}$, respectively, we obtain $w(\sigma^{\prime})=w(\sigma)$. Thus $r({\mathcal{N}})=r({\mathcal{N}}_{m+1})\geq w(\sigma^{\prime})=w(\sigma)$ and so (CP) holds for ${\mathcal{N}}$.
Claim 1: We can and will assume that ${\mathcal{N}}$ is pendantless.
@@
Proof of Claim 1: Suppose ${\mathcal{N}}$ is not pendantless. Then ${\mathcal{N}}$ contains a cherry. Thus by Lemma 6.2 there must be some $x\in X$ such that ${\mathcal{N}}-x$ is a phylogenetic network and (a) or (b) is satisfied for $x$. But this contradicts our assumption made in Claim 0.
@@
Claim 2: We can and will assume that $|L({\mathcal{B}})|\geq 2$.
Proof of Claim 2: If $|L({\mathcal{B}})|\leq 1$, then by applying Lemma 5.6 to ${\mathcal{N}}$ and using induction it follows that (CP) holds. ∎
Claim 3: We can and will assume that no leaf in $L({\mathcal{B}})$ is a component in ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$.
Proof of Claim 3: Suppose that some $x\in L({\mathcal{B}})$ is a component of ${\mathcal{F}}$. Then by Claim 0, the pseudo-network ${\mathcal{N}}-x$ is not a phylogenetic network. So we can apply a (C3)-reduction to $x$ and, since $|L({\mathcal{B}})|\geq 2$ by Claim 2, we can apply Lemma 5.4 to ${\mathcal{F}}$, ${\mathcal{F}}^{\prime}$, ${\mathcal{N}}$ and $x$ to see that (CP) must hold for ${\mathcal{N}}$ by induction. ∎
@@
Claim 4: We can and will assume that every pendant edge incident to ${\mathcal{B}}$ and every edge in ${\mathcal{B}}$ is contained in ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ (or both).
Proof of Claim 4: If a pendant edge were not in ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, then either ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$ would contain an element in $L({\mathcal{B}})$ that is a component in ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$, which contradicts our assumption in Claim 3. And, if there is some $e\in E({\mathcal{B}})$ such that $e$ is neither contained in ${\mathcal{N}}[{\mathcal{F}}]$ nor in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, then since $|L({\mathcal{B}})|\geq 2$ by Claim 2, Lemma 5.5(a) implies that there exists a phylogenetic network ${\mathcal{N}}^{\prime}$ on $X$ that displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ and $r({\mathcal{N}})>r({\mathcal{N}}^{\prime})$. We can then apply induction to see that (CP) holds for ${\mathcal{N}}$.
@@
Proof of Claim 5: Since ${\mathcal{N}}$ is pendantless by Claim 1, $u\neq v$ and $\{x,u\}$ and $\{y,v\}$ are pendant edges of ${\mathcal{B}}$. Moreover, $u$ and $v$ are clearly contained in ${\mathcal{N}}[\gamma]$, and ${\mathcal{N}}[\gamma]$ must contain at least three edges since $u\neq v$. And every edge in ${\mathcal{N}}[\gamma]$ not equal to $\{x,u\}$ and $\{y,v\}$ is contained in $E({\mathcal{B}})$ since ${\mathcal{B}}$ is a pendant blob in ${\mathcal{N}}$. ∎
Claim 6: If $\mathcal{T}$ is any tree in ${\mathcal{F}}$ or in ${\mathcal{F}}^{\prime}$ that contains a cherry $(x,y)$ with $x,y\in L({\mathcal{B}})$, then we can and will assume that out of the four cases (i)–(iv) in Observation 3.1 only case (ii) holds for $x$ and $y$.
@@
Last, suppose Observation 3.1(i) holds, i.e., that $(x,y)$ is a cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. Note that by Claim 0, neither the pseudo-network ${\mathcal{N}}-x$ nor the pseudo-network ${\mathcal{N}}-y$ is a phylogenetic network. Moreover, every edge in ${\mathcal{N}}[\gamma]$ must be contained in both ${\mathcal{N}}[{\mathcal{F}}]$ and ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, otherwise we could use Lemma 5.5(b) and a (C2)(b)(i)-reduction and induction to see that (CP) must hold. Let $e$ and $e^{\prime}$ be the edges in $E({\mathcal{B}})$ incident to $u$ and $v$, respectively, but not in ${\mathcal{N}}[\gamma]$, so that $e\neq e^{\prime}$ and neither $e$ nor $e^{\prime}$ is a cut-edge of ${\mathcal{N}}$. Note that since every edge in ${\mathcal{B}}$ is contained in ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ by Claim 4, and $(x,y)$ is a cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, it follows without loss of generality, that $e$ is in ${\mathcal{N}}[{\mathcal{F}}]$ but not in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, and that $e^{\prime}$ is in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ but not in ${\mathcal{N}}[{\mathcal{F}}]$.
@@
We now show that the length of the path ${\mathcal{N}}[\gamma]$ must be three (noting that we are assuming that it has length at least three). Suppose not. First suppose that the length of ${\mathcal{N}}[\gamma]$ is four. Let $w$ be the vertex in ${\mathcal{N}}[\gamma]$ that is distance two from both $x$ and $y$. Then as $(x,y)$ is a cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, by Claim 4, $w$ must be in the cut-edge that is incident to ${\mathcal{B}}$. Now, if $e$ and $e^{\prime}$ have no vertex in common then, since ${\mathcal{N}}-x$ is a pseudo-network that is not a phylogenetic network, when we remove the edge $\{u,x\}$ from ${\mathcal{N}}$ and suppress the vertex $u$, we must obtain a multi-edge that contains the two vertices in ${\mathcal{B}}$ adjacent to $u$ that are not equal to $x$. But this is impossible since one of these two vertices must be equal to $w$ which would imply that the degree of $w$ in ${\mathcal{N}}$ is 4. And, if $e$ and $e^{\prime}$ have a vertex in common, say $w^{\prime}$, then ${\mathcal{B}}$ must be a 4-cycle, with leaves $x,y$ and another leaf $z$, $z\in X$, such that $\{w^{\prime},z\}$ is a pendant edge of ${\mathcal{B}}$. But this is impossible by Claim 0 since $(z,y)$ is then a cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ and ${\mathcal{N}}-z$ is a phylogenetic network. Finally, note that since $(x,y)$ is cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ and ${\mathcal{B}}$ is a pendant blob, it follows that the length of ${\mathcal{N}}[\gamma]$ cannot be greater than four by Claim 4.
@@
Now, assume that the length of the path ${\mathcal{N}}[\gamma]$ is three. If $e$ and $e^{\prime}$ have no vertex in common, then since ${\mathcal{N}}-x$ is a pseudo-network that is not a phylogenetic network, when we remove the edge $\{u,x\}$ from ${\mathcal{N}}$ and suppress the vertex $u$, we must obtain a multi-edge that contains the two vertices in ${\mathcal{B}}$ adjacent to $u$ that are not equal to $x$. But this is impossible since one of these two vertices must be equal to $v$ as ${\mathcal{N}}[\gamma]$ has length three, and $e$ and $e^{\prime}$ have no vertex in common. And, finally, if $e$ and $e^{\prime}$ have a vertex in common, say $w$, then ${\mathcal{B}}$ is the 3-cycle $u,v,w$ and there is a vertex $w^{\prime}\not=u,v$ with $w^{\prime}\not\in X$ such that $\{w,w^{\prime}\}$ is a non-pendant cut-edge of ${\mathcal{N}}$. So we can apply a (C1)-reduction to $x$, make a new network ${\mathcal{N}}^{\prime}$ on $X\setminus\{x\}$ with $r({\mathcal{N}})>r({\mathcal{N}}^{\prime})$ by removing $x$ and all of the vertices in ${\mathcal{B}}$ and the edges in ${\mathcal{N}}$ containing them and adding the edge $\{w^{\prime},y\}$, and then apply induction to see that (CP) holds.

### v5

@@ 6 Bounding the weight of a cherry picking sequence from above by the hybrid number
Now, in case $r({\mathcal{N}})>0$, we essentially repeatedly look for non cut-edges $e$ in ${\mathcal{N}}$ that are contained in the image of one of the forests displayed by ${\mathcal{N}}$ but not in the other. If we can find such an edge $e$, by carefully considering how the images of the forests are embedded in ${\mathcal{N}}$, we can then complete the proof by removing $e$ from ${\mathcal{N}}$, applying Lemmas 5.4 and 5.6 and some (C2)-reduction, and using the fact that $r({\mathcal{N}})>r({\mathcal{N}}-e)$. We use this tool within a series of claims to show that either we can prove the theorem by induction, or that we can further restrict the way that the forests are embedded in ${\mathcal{N}}$. This eventually allows us to finally apply our induction hypothesis one final time and thus complete the proof.
More specifically, we first show that we can assume that ${\mathcal{N}}$ satisfies a certain property (P). Then, after noting that ${\mathcal{N}}$ must be pendantless (Claim 1), we show that we can also assume that ${\mathcal{N}}$ must have a pendant blob ${\mathcal{B}}$ with at least two leaves such that no leaf of ${\mathcal{B}}$ is a component of ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$ and every edge in ${\mathcal{B}}$ is in ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ (Claims 2, 3 and 4). In Claims 5 and 6 we then show that either ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$ must contain a cherry $(x,y)$ that, out of the four possibilities (i)–(iv) in Observation 3.1, only satisfies (ii). Then in Claims 7–11 we show that we can assume that there is a component in ${\mathcal{F}}$ that contains a cherry $(x,y)$ and a component in ${\mathcal{F}}^{\prime}$ that contains a cherry $(x,z)$ with $x,y,z\in L({\mathcal{B}})$. This allows us to assume that $((x,y),p)$ or $((x,y),(p,q))$ is a pendant subtree of one of the forests for some $x,y,p,q\in L({\mathcal{B}})$ distinct. This permits us to carry out a case analysis in Claims 12–14 in order to complete the proof.
@@
Base Case: Suppose $r({\mathcal{N}})=0$. Then ${\mathcal{N}}$ is a phylogenetic tree on $X$. If $|X|=1$, then $X=\{x\}$, for some $x$, and $(x)$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with $w((x))=0$. Now, suppose $m=|X|>1$. Order the elements in $X$ as $x_{1},x_{2},\dots,x_{m}$ so that, for all $11\leq i\leq m-$ 1, $x_{i}$ is contained in a cherry of ${\mathcal{N}}|{\{x_{i},\dots,x_{m}\}}$ which is possible since any phylogenetic tree with two or more leaves always contains a cherry. Then, by applying Lemma 6.2 and using either a (C1) or (C3) reduction for each $x_{i}$, $11\leq i\leq m-$ 1, depending on whether (a) or (b) holds in Lemma 6.2 for $x_{i}$, respectively, it follows that $(x_{1},\dots,x_{m})$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with $r({\mathcal{N}})=w((x_{1},\dots,x_{m}))=0$. Hence the base case holds.
@@
Note first that we will assume from now on that $|X|\geq 3$ since for $|X|\in\{1,2\}$, we can clearly find a cherry picking sequence $\sigma$ such that $w(\sigma)=0$ and so $r({\mathcal{N}})>0=w(\sigma)$ and hence (CP) holds.
@@
Starting with ${\mathcal{N}}$, repeatedly look for elements $x\in X$ that satisfy (a) there is a cherry $(x,y)$ contained in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, $y\in X\setminus\{x\}$, or (b) $x$ is a component in either ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$, and such that removing $x$ and $e_{x}$ and suppressing the resulting degree-2 vertex results in a phylogenetic network until this is no longer possible. Then we obtain a sequence $(x_{1},\dots,x_{m})$, $x_{i}\in X$, $m\geq 1$, where, for ${\mathcal{N}}_{1}={\mathcal{N}}$, ${\mathcal{N}}_{i+1}={\mathcal{N}}_{i}-x_{i}$, ${\mathcal{F}}[1]={\mathcal{F}}$, ${\mathcal{F}}[i+1]={\mathcal{F}}[i]-x_{i}$, ${\mathcal{F}}^{\prime}[1]={\mathcal{F}}^{\prime}$ and ${\mathcal{F}}^{\prime}[i+1]={\mathcal{F}}^{\prime}[i]-x_{i}$, $1\leq i\leq m$, we have for all $1\leq i\leq m$ that $x_{i}$ satisfies (a) or (b) and ${\mathcal{N}}_{i+1}={\mathcal{N}}_{i}-x_{i}$ is a phylogenetic network on $X\setminus\{x_{1},\dots,x_{i}\}$, ${\mathcal{N}}_{i}$ displays the forests ${\mathcal{F}}[i]$, ${\mathcal{F}}^{\prime}[i]$ and, by Lemma 6.2, $r({\mathcal{N}}_{i})=r({\mathcal{N}})$.
In particular, ${\mathcal{N}}_{m+1}$ is a phylogenetic network on $X\setminus\{x_{1},\dots,x_{m}\}$ with $r({\mathcal{N}}_{{\color[rgb]{0,0,0}m+1}})=q$ that displays ${\mathcal{F}}[m+1]$ and ${\mathcal{F}}^{\prime}[m+1]$, and there is no $x\in X\setminus\{x_{1},\dots,x_{m}\}$ such that ${\mathcal{N}}_{m+1}-x$ is a phylogenetic network and either (a) or (b) holds.
@@
Now suppose that we could show that (CP) holds for ${\mathcal{N}}_{m+1}$ along with the forests ${\mathcal{F}}[m+1]$ and ${\mathcal{F}}^{\prime}[m+1]$ (which we shall do below). Then there must exist a cherry picking sequence $\sigma^{\prime}$ for ${\mathcal{F}}[m+1]$ and ${\mathcal{F}}^{\prime}[m+1]$ with $r({\mathcal{N}}_{m+1})\geq w(\sigma^{\prime})$. Hence, $\sigma=(x_{1},\dots,x_{m})\circ\sigma^{\prime}$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. Since we have applied a (C1) or (C3) reduction to each $x_{i}$ depending on whether (a) or (b) holds for $x_{i}$, respectively, we obtain $w(\sigma^{\prime})=w(\sigma)$. Thus $r({\mathcal{N}})=r({\mathcal{N}}_{m+1})\geq w(\sigma^{\prime})=w(\sigma)$ and so (CP) holds for ${\mathcal{N}}$.
Thus, to show that (CP) holds for ${\mathcal{N}}$, we assume from now on that ${\mathcal{N}}$ satisfies the following property:
(P) There is no $x\in X$ such that ${\mathcal{N}}-x$ is a phylogenetic network on $X\setminus\{x\}$ and either (a) there is a cherry $(x,y)$ contained in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, $y\in X\setminus\{x\}$, or (b) $x$ is a component in either ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$.
Claim 1: ${\mathcal{N}}$ is pendantless.
@@
Proof of Claim 1: Suppose ${\mathcal{N}}$ is not pendantless. Then ${\mathcal{N}}$ contains a cherry. Thus by Lemma 6.2 there must be some $x\in X$ such that ${\mathcal{N}}-x$ is a phylogenetic network and (a) or (b) is satisfied for $x$. But this contradicts the fact that ${\mathcal{N}}$ satisfies property (P). $\blacksquare$
@@
Claim 2: $|L({\mathcal{B}})|\geq 2$.
Proof of Claim 2: If $|L({\mathcal{B}})|\leq 1$, then by applying Lemma 5.8 to ${\mathcal{N}}$ and using induction it follows that (CP) holds. $\blacksquare$
Claim 3: No leaf in $L({\mathcal{B}})$ is a component in ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$.
Proof of Claim 3: Suppose that some $x\in L({\mathcal{B}})$ is a component of ${\mathcal{F}}$. Then by property (P), the pseudo-network ${\mathcal{N}}-x$ is not a phylogenetic network. So we can apply a (C3)-reduction to $x$ and, since $|L({\mathcal{B}})|\geq 2$ by Claim 2, we can also apply Lemma 5.4 to ${\mathcal{F}}$, ${\mathcal{F}}^{\prime}$, ${\mathcal{N}}$ and $x$ to see that (CP) must hold for ${\mathcal{N}}$ by induction. $\blacksquare$
@@
Claim 4: Every pendant edge incident to ${\mathcal{B}}$ and every edge in ${\mathcal{B}}$ is contained in ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ (or both).
Proof of Claim 4: If a pendant edge were not in ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, then either ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$ would contain an element in $L({\mathcal{B}})$ that is a component in ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$, which contradicts our assumption in Claim 3. And, if there is some $e\in E({\mathcal{B}})$ such that $e$ is neither contained in ${\mathcal{N}}[{\mathcal{F}}]$ nor in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, then since $|L({\mathcal{B}})|\geq 2$ by Claim 2, Lemma 5.6(a) implies that there exists a phylogenetic network ${\mathcal{N}}^{\prime}$ on $X$ that displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ and $r({\mathcal{N}})>r({\mathcal{N}}^{\prime})$. We can then apply induction to see that (CP) holds for ${\mathcal{N}}$. $\blacksquare$
@@
Proof of Claim 5: Since ${\mathcal{N}}$ is pendantless by Claim 1, $u\neq v$ and $\{x,u\}$ and $\{y,v\}$ are pendant edges of ${\mathcal{B}}$. Moreover, $u$ and $v$ are clearly contained in ${\mathcal{N}}[\gamma]$, and ${\mathcal{N}}[\gamma]$ must contain at least three edges since $u\neq v$. And every edge in ${\mathcal{N}}[\gamma]$ not equal to $\{x,u\}$ and $\{y,v\}$ is contained in $E({\mathcal{B}})$ since ${\mathcal{B}}$ is a pendant blob in ${\mathcal{N}}$. $\blacksquare$
Claim 6: If $\mathcal{T}$ is any tree in ${\mathcal{F}}$ or in ${\mathcal{F}}^{\prime}$ that contains a cherry $(x,y)$ with $x,y\in L({\mathcal{B}})$, then out of the four cases (i)–(iv) in Observation 3.1 only case (ii) holds for $x$ and $y$.
@@
Last, suppose Observation 3.1(i) holds, i.e., that $(x,y)$ is a cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. Note that by property (P), neither the pseudo-network ${\mathcal{N}}-x$ nor the pseudo-network ${\mathcal{N}}-y$ is a phylogenetic network. Moreover, every edge in ${\mathcal{N}}[\gamma]$ must be contained in both ${\mathcal{N}}[{\mathcal{F}}]$ and ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, otherwise we could use Lemma 5.6(b) and a (C2)(b)(i)-reduction and induction to see that (CP) must hold. Let $e$ and $e^{\prime}$ be the edges in $E({\mathcal{B}})$ incident to $u$ and $v$, respectively, but not in ${\mathcal{N}}[\gamma]$, so that $e\neq e^{\prime}$ and neither $e$ nor $e^{\prime}$ is a cut-edge of ${\mathcal{N}}$. Note that since every edge in ${\mathcal{B}}$ is contained in ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ by Claim 4, and $(x,y)$ is a cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, it follows without loss of generality, that $e$ is in ${\mathcal{N}}[{\mathcal{F}}]$ but not in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, and that $e^{\prime}$ is in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ but not in ${\mathcal{N}}[{\mathcal{F}}]$.
@@
We now show that the length of the path ${\mathcal{N}}[\gamma]$ must be three (noting that we are assuming that it has length at least three). Suppose not. First suppose that the length of ${\mathcal{N}}[\gamma]$ is four. Let $w$ be the vertex in ${\mathcal{N}}[\gamma]$ that is distance two from both $x$ and $y$. Then, by Claim 4, $e$ and $e^{\prime}$ must be in either ${\mathcal{N}}[{\mathcal{F}}]$ or ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$. Since $(x,y)$ is a cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, it follows that the edge incident to ${\mathcal{B}}$ and not in ${\mathcal{N}}[\gamma]$ is neither in ${\mathcal{N}}[{\mathcal{F}}]$ nor ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$. Hence, $w$ must be contained in the cut-edge of ${\mathcal{N}}$ that is incident to ${\mathcal{B}}$. Now, if $e$ and $e^{\prime}$ have no vertex in common then, since ${\mathcal{N}}-x$ is a pseudo-network that is not a phylogenetic network, when we remove the edge $\{u,x\}$ from ${\mathcal{N}}$ and suppress the vertex $u$, we must obtain a multi-edge that contains the two vertices in ${\mathcal{B}}$ adjacent to $u$ that are not equal to $x$. But this is impossible since one of these two vertices must be equal to $w$, and the degree of $w$ in ${\mathcal{N}}$ is 3. And, if $e$ and $e^{\prime}$ have a vertex in common, say $w^{\prime}$, then ${\mathcal{B}}$ must be a 4-cycle, with leaves $x,y$ and another leaf $z$, $z\in X$, such that $\{w^{\prime},z\}$ is a pendant edge of ${\mathcal{B}}$. But this is impossible by property (P) since $(z,y)$ is then a cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ and ${\mathcal{N}}-z$ is a phylogenetic network. Finally, note that since $(x,y)$ is cherry in both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ and ${\mathcal{B}}$ is a pendant blob, it follows that the length of ${\mathcal{N}}[\gamma]$ cannot be greater than four by Claim 4.
@@
Now, assume that the length of the path ${\mathcal{N}}[\gamma]$ is three. If $e$ and $e^{\prime}$ have no vertex in common, then since ${\mathcal{N}}-x$ is a pseudo-network that is not a phylogenetic network, when we remove the edge $\{u,x\}$ from ${\mathcal{N}}$ and suppress the vertex $u$, we must obtain a multi-edge that contains the two vertices in ${\mathcal{B}}$ adjacent to $u$ that are not equal to $x$. But this is impossible since one of these two vertices must be equal to $v$ as ${\mathcal{N}}[\gamma]$ has length three, and $e$ and $e^{\prime}$ have no vertex in common. And, finally for the case where ${\mathcal{N}}[\gamma]$ has length three, if $e$ and $e^{\prime}$ have a vertex in common, say $c$, then ${\mathcal{B}}$ is the 3-cycle $u,v,c$ and there is a vertex $c^{\prime}\not=u,v$ with $c^{\prime}\not\in X$ such that $\{c,c^{\prime}\}$ is a non-pendant cut-edge of ${\mathcal{N}}$. So we can apply a (C1)-reduction to $x$, make a new network ${\mathcal{N}}^{\prime}$ on $X\setminus\{x\}$ with $r({\mathcal{N}})>r({\mathcal{N}}^{\prime})$ by removing $x$ and all of the vertices in ${\mathcal{B}}$ and the edges in ${\mathcal{N}}$ containing them and adding the edge $\{c^{\prime},y\}$, and then apply induction to see that (CP) holds. $\blacksquare$